a girl has 24 quarters and dimes. togther they had up to 3.60 how many quarters does she have?
Found 2 solutions by elima, checkley71:
Answer by elima(1433)

a girl has 24 quarters and dimes. togther they had up to 3.60 how many quarters does she have?
q+d=24
.25q+.1d=3.60
=====================
d=24-q
.25q+.1(24-q)=3.60
.25q+2.4-.1q=3.60
.15q=3.6-2.4
.15q=2.4
q=8
================
8+d=24
d=24-8
d=16
:)

Answer by checkley71(8403)

Q+D=24 OR D=24-Q
.25Q+.10(24-Q)=3.60
.25Q+2.4-.10Q=3.60
.15Q=3.60-2.40
.15Q=1.20
Q=1.20/.15
Q=8 ANSWER FOR THE NUMBER OF QUARTERS.
24-8=16 NUMBER OF DIMES.
PROOF
.25*8+.10*16=3.60
2.00+1.60=3.60
3.60=3.60